OpenStudy (anonymous):

A 60 ounce glass bottle of juice sells for $4.20 while a 1.5 liter of the same juice sells in a plastic bottle for $3.06. Compare the unit prices of the juice for the two containers. (1 liter is approximately 34 ounces) a. The juice in the glass bottle is 1 cent cheaper per ounce. b. The juice in the plastic bottle is 1 cent cheaper per ounce. c. The juice in the glass bottle is 10 cents cheaper per ounce. d. The juice in the plastic bottle is 10 cents cheaper per ounce.

OpenStudy (anonymous):

The answer is D

OpenStudy (johnweldon1993):

Okay so lets put it in terms of ounces... 1 liter is roughly 34 ounces...and we have a 1.5 liter...so that is 34 + half of 34 which is 17...so 34 + 17 = 51 ounces... Glass: 60 ounces for 4.20 Plastic: 51 ounces for 3.06 if we want to find out how much per 1 ounces...we divide the 4.20 by 60...and we divide 51 by 3.06

OpenStudy (anonymous):

D?

OpenStudy (johnweldon1993):

@ashleyhelget does that make sense?

OpenStudy (anonymous):

is it d

OpenStudy (mertsj):

60 ounces for 4.20 means 60 ounces for 420 cents which would be : \[\frac{420 cents}{60 ounces}=7 cents/ounce\]

OpenStudy (mertsj):

1.5 liters = 1.5 time 34 ounces = 51 ounces. That would be 306 cents for 51 ounces or: \[\frac{306 cents}{51 ounces}= 6 cents/ounce\]

OpenStudy (mertsj):

Now you choose the right answer. Which one is it?

OpenStudy (texaschic101):

did you get it yet ?

Directrix (directrix):

@MathDude000 Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct